OpenStudy (anonymous):

is the system consistent or inconsistent? 2x+3y=5 and 4x+6y=19

OpenStudy (anonymous):

What do you see if you double the first equation?

OpenStudy (anonymous):

4x,6y, and25

OpenStudy (whpalmer4):

doubling 5 gives you 25?

OpenStudy (anonymous):

I know im confused

OpenStudy (whpalmer4):

\[2x+3y=5\]\[2*2x+2*3y=2*5\]\[4x + 6y = 10\]

OpenStudy (anonymous):

im not confused now thank you